Required information

[The following information applies to the questions displayed below.]

At the beginning of Year 1, Oak Consulting had the following normal balances in its accounts:

Account Balance
Cash $ 29,200
Accounts receivable 16,400
Accounts payable 12,200
Common stock 20,200
Retained earnings 13,200


The following events apply to Oak Consulting for Year 1:

  1. Provided $74,500 of services on account.
  2. Incurred $3,000 of operating expenses on account.
  3. Collected $50,000 of accounts receivable.
  4. Paid $36,000 cash for salaries expense.
  5. Paid $13,680 cash as a partial payment on accounts payable.
  6. Paid a $8,600 cash dividend to the stockholders.

d-1. Record the closing entries in the general journal. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

Record the closure of service revenue.

Record the closure of expenses account.

Record the payment of dividends.

d-2. What is the amount of net income for the year?

Net income
Answer #1

d-1) Closing entry

No General Journal Debit Credit
1 Service revenue 74500
Retained earnings 74500
(To close revenue)
2 Retained earnings 39000
Operating expense 3000
Salaries expense 36000
(To close expense)
3 Retained earnings 8600
Dividend 8600
(To close dividend)

d-2) Net income = Revenue-Expense = 74500-39000 = 35500